OpenStudy (anonymous):

using distributive law calculate the value of 2-3(4-6)+7=

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

what are your options?

OpenStudy (skullpatrol):

$$\Huge a-b=a+(-b)$$ $$\Huge 2-3(4-6)+7$$$$\Huge= 2+(-3)(4-6)+7 $$

OpenStudy (anonymous):

I got 15 😌

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

distributive law is showing how multiplication affects an addition or subtraction ex: a(b+c)=AxB+AxC

OpenStudy (anonymous):

can u break it down simpler :(

OpenStudy (skullpatrol):